OpenStudy (anonymous):

Simplify ((27y^4)/(25x^13z16))^(1/3)

OpenStudy (anonymous):

Is that 16z or z16?!?

OpenStudy (anonymous):

Sorry \[\sqrt[3]{\frac{27y^4}{25x^13z^{16}}}\]

OpenStudy (anonymous):

This is what I got but I'm not sure if this is right. You might wanna check it.. \[(3y^3/z^5)\sqrt[3]{y/(25x3z)}\]

OpenStudy (anonymous):

It's not simplified... I got \[\frac{3y \sqrt[3]{5}}{5\sqrt[3]{x^{13}}\sqrt[3]{z^{16}}}\]

OpenStudy (anonymous):

How did you take out 5? Cube root of 25 is not 5.. I'm confuse on how you did it..

OpenStudy (anonymous):

Because to take the radical out of the denominator I multipled numerator and denominator by the cube root of 5x^2z^2